OpenStudy (anonymous):

Bo and Erica are yoga instructors. Between the two of them, they teach 33 yoga classes each week. If Erica teaches 15 fewer than twice as many as Bo, how many classes does each instructor teach per week?

jimthompson5910 (jim_thompson5910):

B = # of classes Bo teaches E = # of classes Erica teaches B+E = 33 (they teach 33 classes total) E = 2B - 15 (Erica teaches 15 fewer than twice as many as Bo)

jimthompson5910 (jim_thompson5910):

B+E = 33 B+2B - 15 = 33 ... plug in E = 2B - 15 3B - 15 = 33 3B = 33 + 15 3B = 48 B = ???

OpenStudy (anonymous):

16?

jimthompson5910 (jim_thompson5910):

B = 16 E = 2B - 15 E = 2*16 - 15 E = 32 - 15 E = 17 So Bo teaches 16 classes per week while Erica teaches 17 classes per week

OpenStudy (anonymous):

im so sorry i dont get it

OpenStudy (anonymous):

4x+y=12 2x-2y=6

jimthompson5910 (jim_thompson5910):

solve 4x+y=12 for y then plug it into the second equation solve that for x and use it to find y